# Ambient Noise Full Waveform Inversion with Neural Operators

### v3

@@ 5 Conclusions
The node data can be accessed from the IRIS Data Management Center (BASIN, 2018). The network codes are 4M and 6J. The code, models, and processed data are available at https://github.com/caifeng-zou/ANFWI_HNO (Zou, 2025).
@@ Acknowledgements
ZER acknowledges financial support from the David and Lucile Packard Foundation. CZ and RWC acknowledge support from NSF EAR-2105358 and EAR-2438773. FCL acknowledges financial support from the Statewide California Earthquake Center (SCEC) based on Award Number DE-SC0016520 from the U.S. Department of Energy and NSF EAR-2438772.
@@
Figure S1: Loss curves. The batch size is 256256256. We use an Adam optimizer with a learning rate of 0.0010.0010.001 and a scheduler that decays the learning rate by half every 303030 epochs.
@@
Velocity model Covariance $var$ $len\_scale$ $nu$ Background
$V_{S}=BG_{S}\times(1+RF_{S}/100)$ $RF_{S}$: Matérn 400 (64, 8) 1.5 $BG_{S}$: 1D from Lee et al. (2014)
$V_{P}=BG_{P}\times(1+RF_{P}/100)$ $RF_{P}$: Matérn 4 (64, 8) 2.5 $BG_{P}(V_{S})$: Brocher (2005)

### v4

@@ 6 Conclusions
The node data can be accessed from the IRIS Data Management Center (BASIN2018). The network codes are 4M and 6J. The code, models, and processed data are available at https://github.com/caifeng-zou/ANFWI_HNO (caifeng_zou_2025_15061657).
@@ Acknowledgements
ZER acknowledges financial support from the David and Lucile Packard Foundation. CZ and RWC acknowledge support from NSF EAR-2105358 and EAR-2438773. FCL acknowledges financial support from the Statewide California Earthquake Center (SCEC) based on Award Number DE-SC0016520 from the U.S. Department of Energy and NSF EAR-2438772. We sincerely thank Fenglin Niu as the editor, Tariq Alkhalifah, and an anonymous reviewer for their careful reading and constructive feedback, which led to substantial improvements in this paper. We thank all the volunteers who helped deploy the nodal arrays and Patricia Persaud for coordinating the BASIN project.
@@
Figure S1: Loss curves. The batch size is $256$. We use an Adam optimizer with a learning rate of $0.001$ and a scheduler that decays the learning rate by half every $30$ epochs.
@@
Figure S3: Synthetic FWI results for the velocity model from li2023shear.
@@
Figure S5: Overall MSE (mean squared error) in the FWI process for SB1 with different initial models, corresponding to Figure 8. Higher frequency data is progressively incorporated, as in synthetic tests. We use an Adam optimizer with a learning rate of $0.05$ and a scheduler that decays the learning rate by half every $5$ epochs. The batch size is $64$.
@@
Velocity model Covariance varvar len_scalelen\_scale nunu Background
$V_{S}=BG_{S}\times(1+RF_{S}/100)$ $RF_{S}$: Matérn 400 (64, 8) 1.5 $BG_{S}$: 1D from lee2014full
$V_{P}=BG_{P}\times(1+RF_{P}/100)$ $RF_{P}$: Matérn 4 (64, 8) 2.5 $BG_{P}(V_{S})$: brocher2005empirical